# SADA: Safe and Adaptive Inference with Multiple Black-Box Predictions

### v1

@@ 2 Problem Setup
Table 1: Data structure in SSL with multiple sets of predicted labels.
@@
Overarching goal
The overarching goal of this paper is to propose a safe and adaptive algorithm that constructs valid and more informative inference for $\bm{\theta}$$^{*}$ by leveraging multiple predictions, regardless of their individual quality. Importantly, we do not require the predictions $\widehat{Y}_{k}$ to share the same magnitude or form either with each other or with $Y$. For example, when $Y$ is a binary label, each prediction $\widehat{Y}_{k}$ may be either categorical or continuous. We allow the generation process of $\widehat{Y}_{k}$ to be a black box, potentially depending not only on the observed feature $\mathbf{X}$ but also on some unobservable or latent variables.
3 Key Idea: Illustration with Mean Estimation
@@
We elaborate the intuition behind our approach using the example of mean estimation. Consider $\mathbf{s}(\mathbf{x},y;\theta)=y-\theta$, which corresponds to the estimand $\theta^{*}=\mathbb{E}(Y)$, the outcome mean. The naive estimator is the averaged outcome of the labeled samples, $\widehat{\theta}^{\text{nv}}=n^{-1}\sum_{i=1}^{n}y_{i}$. While this estimator is unbiased, it may suffer from high variance and mean squared error due to the limited size of the labeled dataset. To improve efficiency by leveraging the $K$ machine learning-predicted outcomes $\widehat{\mathbf{y}}_{i}$ from both labeled and unlabeled data, we introduce a family of unbiased estimators indexed by weights $\bm{\omega}=(\omega_{1},\dots,\omega_{K})^{\mathrm{\scriptscriptstyle T}}$:
@@
This family covers several existing methods as special cases. For example, when $\bm{\omega}=\mathbf{0}$, it reduced to the naive estimator $\widehat{\theta}^{\text{nv}}$. When there is only one prediction, i.e. $K=1$, and $\omega=1$, it turns out to be the PPI estimator (Angelopoulos et al.,, 2023):
@@
$\mathbb{E}[\{\widehat{\theta}(\bm{\omega})-\theta^{*}\}^{2}]=\underbrace{\frac{1}{n}\operatorname{var}(Y)}_{\text{naive estimator}}+\underbrace{\frac{N}{n(N-n)}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\operatorname{var}(\widehat{\mathbf{Y}})\bm{\omega}-\frac{2}{n}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\operatorname{cov}(\widehat{\mathbf{Y}},Y)}_{\text{additional term}}.$ (3)
@@
where $\overline{y}=n^{-1}\sum_{i=1}^{n}y_{i}$, and $\overline{\widehat{\mathbf{y}}}=N^{-1}\sum_{i=1}^{N}\widehat{\mathbf{y}}_{i}$.
In this paper, we refer to the resulting optimal estimator $\widehat{\theta}(\widehat{\bm{\omega}}^{\text{opt}})$ as the proposed SADA estimator, denoted as $\widehat{\theta}^{\text{sada}}$. In what follows, for certain theoretical analyses and illustrations, we do not distinguish between the estimator $\widehat{\theta}(\bm{\omega}^{\text{opt}})$ and the SADA estimator $\widehat{\theta}^{\text{sada}}$, as they are asymptotically equivalent. This simplification is made when it does not lead to confusion.
The SADA estimator $\widehat{\theta}^{\text{sada}}$ enjoys the following two attractive properties.
Safety
@@
$\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}\}=\underbrace{\frac{1}{n}\operatorname{var}(Y)}_{\text{naive estimator}}-\underbrace{\left(\frac{1}{n}-\frac{1}{N}\right)\operatorname{cov}(\widehat{\mathbf{Y}},Y)^{\mathrm{\scriptscriptstyle T}}\operatorname{var}(\widehat{\mathbf{Y}})^{-1}\operatorname{cov}(\widehat{\mathbf{Y}},Y)}_{\text{efficiency gain}}.$ (5)
The efficiency gain in (5) is always non-negative regardless of the quality of the predictions. It vanishes to zero if and only if $\operatorname{cov}(\widehat{\mathbf{Y}},Y)=\mathbf{0}$, that means, none of the predictions are correlated with the ground truth $Y$. In this worst-case scenario, the optimal weight is automatically assigned to zero (i.e., $\bm{\omega}=\mathbf{0}$), reducing to the naive estimator. Except this worst-case scenario, the SADA estimator offers a guaranteed positive efficiency gain over the naive method and enables more informative inference for $\theta^{*}$ (see Theorem 1 and Appendix B for details in the general case).
Adaptivity
@@
First, consider the case where one of the predictions perfectly matches the ground truth, for example, $\widehat{Y}_{1}\equiv Y$. As shown in Appendix C, the optimal weight in this case is $\bm{\omega}^{\text{opt}}=(1,0,\dots,0)^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. This means that the algorithm selects the most accurate prediction, $\widehat{Y}_{1}$, for estimation while discarding the less informative ones. Then, the SADA estimator turns out to be $N^{-1}\sum_{i=1}^{N}\widehat{y}_{1,i}$, and $\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}\}=N^{-1}\operatorname{var}(Y)$, same as the oracle estimator who knows the ground truth of the unlabeled data. In this case, the SADA estimator converges at a faster rate of $N^{-1/2}$.
Second, consider the restricted case where the predictions are deterministic functions of the available feature $\mathbf{X}$, i.e., $Y_{k}=\widehat{f}_{k}(\mathbf{X})$. Then the best prediction one can expect to fit the outcome is the conditional mean, $\mathbb{E}(Y\mid\mathbf{X})$, which minimizes the mean squared error between $Y$ and $f(\mathbf{X})$. Assume that $\widehat{Y}_{1}\equiv\mathbb{E}(Y\mid\mathbf{X})$. As shown in Appendix C, the optimal weight in this case also equals $\bm{\omega}^{\text{opt}}=(1,0,\dots,0)^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. We show in the following proposition that $\widehat{\theta}^{\text{sada}}$ achieves the semiparametric efficiency bound (see Appendix A for the definition).
@@
Denote $r_{i}=1$ for labeled units $i=1,\dots,n$ and 0 for unlabeled units. Suppose $n/N\to\pi\in(0,1)$ as $n,N\to\infty$. Assume $\widehat{Y}_{k}=\widehat{f}_{k}(\mathbf{X})$ for $k=1,\dots,K$. Then the efficient influence function (EIF) for estimating $\theta^{*}$, based on labeled data $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ and unlabeled data $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$, is $\phi_{\text{eif}}(r,\mathbf{x},y)=r\pi^{-1}\{y-\mathbb{E}(Y\mid\mathbf{x})\}+\mathbb{E}(Y\mid\mathbf{x})-\theta^{*}$. For the estimator $\widehat{\theta}^{\text{sada}}$, we have $\sqrt{N}(\widehat{\theta}^{\text{sada}}-\theta^{*})\xrightarrow{d}\mathcal{N}\{0,\mathbb{E}(\phi_{\text{eif}}^{2})\}$, which attains the semiparametric efficiency bound.
@@
In case of the mean estimation (or, more generally, a scalar parameter $\theta^{*}$) with $K=1$, the SADA estimator $\widehat{\theta}^{\text{sada}}$ is equivalent to the PPI++ estimator (Angelopoulos et al.,, 2024). As shown later, for a vector-valued parameter, the PPI++ estimator is generally less efficient than the proposed algorithm. Moreover, the PPI++ approach cannot, in general, leverage multiple predictions ($K>1$) simultaneously.
4 Full Protocol: Safe and Adaptive Aggregation of Multiple Predictions
Considering the general parameter $\bm{\theta}$$^{*}$ defined in (1), we construct a family of unbiased estimators, $\widehat{\bm{\theta}}(\mathcal{W})$, indexed by the tuning parameters $\mathcal{W}=(\mathcal{W}_{1}^{\mathrm{\scriptscriptstyle T}},\mathcal{W}_{2}^{\mathrm{\scriptscriptstyle T}},\dots,\mathcal{W}_{K}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{(Kp)\times p}$, which solves
Noting that the second term has zero expectation regardless of the choice of $\mathcal{W}$ or the performance of $\widehat{\mathbf{y}}$, therefore, equation (6) is always a feasible estimating equation for $\bm{\theta}$$^{*}$. In particular, the naive estimator $\widehat{\bm{\theta}}^{\text{nv}}$ is a special case with $\mathcal{W}=\mathbf{0}$. When there is only one prediction $(K=1)$: (i) if $\mathcal{W}=\mathbf{I}$, $\widehat{\bm{\theta}}(\mathcal{W})$ reduces to the PPI estimator (Angelopoulos et al.,, 2023); (ii) if $\mathcal{W}=\omega\mathbf{I}$ with the tuning parameter $\omega$ selected optimally, it reduces to the PPI++ estimator (Angelopoulos et al.,, 2024).
We propose to identify the best estimator by minimizing the mean squared error of $\widehat{\bm{\theta}}(\mathcal{W})$.
Proposition 2.
Among the family of estimators $\widehat{\bm{\theta}}(\mathcal{W})$, the optimal tuning parameter, $\mathcal{W}^{\text{opt}}$, that minimizes the mean squared error loss such that $\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]\preceq\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}]$ for any $\mathcal{W}\in\mathbb{R}^{(Kp)\times p}$ is
The proof of Proposition 2 is given in Appendix D. Proposition 2 implies that $\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})$ performs no worse than the naive estimator regardless of the quality of predictions. To clarify this, one can compute that $\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]=n^{-1}\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{opt}}\mathbf{H}^{-1}$, where
The matrix ${\bm{\Sigma}}_{\text{opt}}$ consists of two terms: the first term corresponds to the variance of the naive estimator, ${\bm{\Sigma}}_{\text{nv}}$, defined in (2), while the second term is a positive semi-definite matrix that represents an efficiency gain. Practically, one can estimate the optimal weight, $\mathcal{W}^{\text{opt}}$, by
where $\widehat{\bm{\theta}}$ is some consistent estimator of $\bm{\theta}$$^{*}$, for example, $\widehat{\bm{\theta}}^{\text{nv}}$. Then, the proposed SADA estimator is $\widehat{\bm{\theta}}^{\text{sada}}=\widehat{\bm{\theta}}(\widehat{\mathcal{W}}^{\text{opt}})$. We briefly summarize the whole protocol for calculating the proposed estimator $\widehat{\theta}^{\text{sada}}$ in Figure 1.
Figure 1: Protocol for computing the proposed SADA estimator $\widehat{\theta}^{\text{sada}}$.
4.1 Properties of the SADA estimator
We impose the following regularity assumptions.

### v2

@@ 2 Setup
Table 1: Data structure in semi-supervised learning with multiple sets of predicted labels.
@@
3 SADA for inference
In this section, we consider the estimation and statistical inference of $\bm{\theta}$$^{*}$ defined in (1). Since the outcome is observed only in the labeled data set $\mathcal{L}$, a naive estimator, $\widehat{\bm{\theta}}^{\text{nv}}$, based solely on the labeled data, solves
Assume that the Hessian matrix $\mathbf{H}=\mathbb{E}\{\partial\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}/\partial{\bm{\theta}}\}$ exists and is nonsingular. Under some regularity conditions (see Assumption 1), $\widehat{\bm{\theta}}^{\text{nv}}$ is consistent and asymptotically normal, and its mean squared error, up to a negligible term, is
where ${\bm{\Sigma}}_{\text{nv}}=\operatorname{var}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}$. Based on (2), valid confidence intervals for $\bm{\theta}$$^{*}$ can be constructed using $\widehat{\bm{\theta}}^{\text{nv}}$. However, since the naive estimator uses only the labeled subset of the sample, it can be substantially less efficient, especially when $n$ is small relative to $N$.
To exploit the unlabeled data, we construct a family of unbiased estimators, $\widehat{\bm{\theta}}(\mathcal{W})$, indexed by the tuning parameter $\mathcal{W}=(\mathcal{W}_{1}^{\mathrm{\scriptscriptstyle T}},\mathcal{W}_{2}^{\mathrm{\scriptscriptstyle T}},\dots,\mathcal{W}_{K}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{(Kp)\times p}$, which solves
Noting that the augmented term has zero expectation regardless of the choice of $\mathcal{W}$ or the performance of $\widehat{\mathbf{y}}$, therefore, equation (3) is always a feasible estimating equation for $\bm{\theta}$$^{*}$. In particular, the naive estimator $\widehat{\bm{\theta}}^{\text{nv}}$ is a special case with $\mathcal{W}=\mathbf{0}$. When there is only one prediction $(K=1)$: (i) if $\mathcal{W}=\mathbf{I}$, $\widehat{\bm{\theta}}(\mathcal{W})$ reduces to the PPI estimator (Angelopoulos et al., 2023); (ii) if $\mathcal{W}=\omega\mathbf{I}$ with the tuning parameter $\omega$ selected optimally, it reduces to the PPI++ estimator (Angelopoulos et al., 2024). However, PPI and PPI++ do not provide a principled way to handle settings with multiple predictions, and even in the special case $K=1$, they do not in general achieve optimality within the family (3) in terms of the smallest possible asymptotic variance. For inference purposes, we therefore propose to identify the best estimator within the family (3) by minimizing the mean squared error of $\widehat{\bm{\theta}}(\mathcal{W})$, as formally stated in the following proposition.
Proposition 1.
Among the family of estimators $\widehat{\bm{\theta}}(\mathcal{W})$ solving (3), the optimal tuning parameter, $\mathcal{W}^{\text{opt}}$, that minimizes the mean squared error loss such that $\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]\preceq\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}]$ for any $\mathcal{W}\in\mathbb{R}^{(Kp)\times p}$ is
The proof of Proposition 1 is given in Supplementary Section S3. Proposition 1 implies that the optimal weight $\mathcal{W}^{\text{opt}}$ exists uniquely and $\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})$ consistently attains asymptotic variance no larger than that of the labeled-only estimator, regardless of the quality of predictions. To clarify this, one can compute that $\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]=n^{-1}\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{opt}}\mathbf{H}^{-1}$, where
The matrix ${\bm{\Sigma}}_{\text{opt}}$ consists of two components: the first corresponds to the variance of the naive estimator, ${\bm{\Sigma}}_{\text{nv}}$, defined in (2), and the second is a positive semi-definite matrix that represents the efficiency gain. We refer to this as a safety property, since for inference purposes, confidence intervals constructed from $\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})$ are always more informative than those based on the naive estimator $\widehat{\bm{\theta}}^{\text{nv}}$. Beyond this safeguard, the optimal weight has another appealing adaptive feature: it assigns higher weights to more accurate predictions and downweights less accurate ones, thereby achieving greater efficiency. In particular, when some prediction is “perfect” in some sense, the procedure automatically exploits it to attain a faster convergence rate or the semiparametric efficiency. We build intuition for how the optimal weight operates and illustrate these properties using mean estimation in Section 3.1, and we formalize them in Section 3.2.
In practice, the optimal weight, $\mathcal{W}^{\text{opt}}$, can be replaced by a consistent estimator, for example,
where $\widehat{\bm{\theta}}$ is a consistent estimator of $\bm{\theta}$$^{*}$ such as $\widehat{\bm{\theta}}^{\text{nv}}$. We denote the resulting estimator $\widehat{\bm{\theta}}^{\text{sada}}=\widehat{\bm{\theta}}(\widehat{\mathcal{W}}^{\text{opt}})$ as the SADA estimator. The whole protocol for calculating $\widehat{\bm{\theta}}^{\text{sada}}$ is summarized in Figure 1.
Figure 1: Protocol for computing the proposed SADA estimator $\widehat{\bm{\theta}}^{\text{sada}}$.
3.1 Illustration with mean estimation
@@
We elaborate the intuition behind our approach using the example of mean estimation. Consider $\mathbf{s}(\mathbf{x},y;\theta)=y-\theta$, which corresponds to the estimand $\theta^{*}=\mathbb{E}(Y)$, the outcome mean. The naive estimator is the averaged outcome of the labeled samples, $\widehat{\theta}^{\text{nv}}=n^{-1}\sum_{i=1}^{n}y_{i}$. To improve efficiency by leveraging $K$ machine learning-predicted outcomes $\widehat{\mathbf{y}}_{i}$ from both labeled and unlabeled data, we consider the family of unbiased estimators $\widehat{\theta}(\bm{\omega})$, indexed by weights $\bm{\omega}=(\omega_{1},\dots,\omega_{K})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{K}$, obtained by solving (3), namely,
@@
As noted previously, when $\bm{\omega}=\mathbf{0}$, it reduced to the naive estimator $\widehat{\theta}^{\text{nv}}$. When there is only one prediction, i.e., $K=1$, and $\omega=1$, it turns out to be the PPI estimator (Angelopoulos et al., 2023):
@@
$\mathbb{E}[\{\widehat{\theta}(\bm{\omega})-\theta^{*}\}^{2}]=\frac{1}{n}\operatorname{var}(Y)+\underbrace{\frac{N}{n(N-n)}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\operatorname{var}(\widehat{\mathbf{Y}})\bm{\omega}-\frac{2}{n}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\operatorname{cov}(\widehat{\mathbf{Y}},Y)}_{\text{additional variance term}}.$ (4)
@@
where $\overline{y}=n^{-1}\sum_{i=1}^{n}y_{i}$, and $\overline{\widehat{\mathbf{y}}}=N^{-1}\sum_{i=1}^{N}\widehat{\mathbf{y}}_{i}$. Then, the SADA estimator is given by $\widehat{\theta}^{\text{sada}}=\widehat{\theta}(\widehat{\bm{\omega}}^{\text{opt}})$. In what follows, for certain theoretical analyses and illustrations, we do not distinguish between the estimator $\widehat{\theta}(\bm{\omega}^{\text{opt}})$ and the SADA estimator $\widehat{\theta}^{\text{sada}}$, as they are asymptotically equivalent. We adopt this simplification when it does not cause confusion. We next demonstrate two key properties of the SADA estimator: safety and adaptivity.
Safety.
@@
$\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}\}=\underbrace{\frac{1}{n}\operatorname{var}(Y)}_{\text{variance of }\widehat{\theta}^{\text{nv}}}-\underbrace{\left(\frac{1}{n}-\frac{1}{N}\right)\operatorname{cov}(\widehat{\mathbf{Y}},Y)^{\mathrm{\scriptscriptstyle T}}\operatorname{var}(\widehat{\mathbf{Y}})^{-1}\operatorname{cov}(\widehat{\mathbf{Y}},Y)}_{\text{efficiency gain}}.$ (6)
The efficiency gain in (6) is always non-negative regardless of the quality of the predictions. It vanishes to zero if and only if $\operatorname{cov}(\widehat{\mathbf{Y}},Y)=\mathbf{0}$; that means, none of the predictions are correlated with the ground truth $Y$. In this worst-case scenario, the optimal weight is automatically assigned to zero (i.e., $\bm{\omega}=\mathbf{0}$), reducing to the naive estimator. Except for this worst-case scenario, the SADA estimator offers a guaranteed positive efficiency gain over the naive method and enables more informative inference for $\theta^{*}$.
Adaptivity.
@@
First, consider the case where one of the predictions perfectly matches the ground truth, for example, $\widehat{Y}_{1}\equiv Y$. As proved in Supplementary Section S4, the optimal weight in this case is $\bm{\omega}^{\text{opt}}=(1,0,\dots,0)^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. This means that the algorithm selects the most accurate prediction, $\widehat{Y}_{1}$, for estimation while discarding the less informative ones. Then, the SADA estimator turns out to be $N^{-1}\sum_{i=1}^{N}\widehat{y}_{1,i}$, and $\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}\}=N^{-1}\operatorname{var}(Y)$, same as the oracle estimator who knows the ground truth of the unlabeled data. In this case, the SADA estimator converges at a faster rate of $N^{-1/2}$.
Second, consider the restricted case where the predictions are deterministic functions of the available feature $\mathbf{X}$, i.e., $Y_{k}=\widehat{f}_{k}(\mathbf{X})$. Then the best prediction one can expect to fit the outcome is the conditional mean, $\mathbb{E}(Y\mid\mathbf{X})$, which minimizes the mean squared error between $Y$ and $f(\mathbf{X})$. Assume that $\widehat{Y}_{1}\equiv\mathbb{E}(Y\mid\mathbf{X})$. As proved in Supplementary Section S4, the optimal weight in this case also equals $\bm{\omega}^{\text{opt}}=(1,0,\dots,0)^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. We show in the following proposition that $\widehat{\theta}^{\text{sada}}$ achieves the semiparametric efficiency bound (see Supplementary Section S1 for the definition).
@@
Denote $r_{i}=1$ for labeled units $i=1,\dots,n$ and 0 for unlabeled units. Suppose $n/N\to\pi\in(0,1)$ as $n,N\to\infty$. Assume $\widehat{Y}_{k}=\widehat{f}_{k}(\mathbf{X})$ for $k=1,\dots,K$. Then, the efficient influence function (EIF) for estimating $\theta^{*}$, based on labeled data $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ and unlabeled data $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$, is $\phi_{\text{eif}}(r,\mathbf{x},y)=r\pi^{-1}\{y-\mathbb{E}(Y\mid\mathbf{x})\}+\mathbb{E}(Y\mid\mathbf{x})-\theta^{*}$. Suppose $\widehat{Y}_{k}\equiv\mathbb{E}(Y\mid\mathbf{X})$ for some $k$, then we have $\sqrt{N}(\widehat{\theta}^{\text{sada}}-\theta^{*})\xrightarrow{d}\mathcal{N}\{0,\mathbb{E}(\phi_{\text{eif}}^{2})\}$, which attains the semiparametric efficiency bound.
@@
In case of the mean estimation (or, more generally, a scalar parameter $\theta^{*}$) with $K=1$, the SADA estimator $\widehat{\theta}^{\text{sada}}$ is conceptually equivalent to the PPI++ estimator (Angelopoulos et al., 2024). However, for a vector-valued parameter, the PPI++ estimator does not in general guarantee a smaller variance than the labeled-only estimator and is thus less efficient than SADA. This phenomenon is evident in our empirical results in Section 6. In addition, the PPI++ approach cannot simultaneously leverage multiple predictions when $K>1$.
3.2 Theoretical guarantees
In this section, we formally establish the theoretical properties of the SADA estimator $\widehat{\bm{\theta}}^{\text{sada}}$. We impose the following regularity conditions that ensure identification, smoothness, and well-behaved large-sample properties of the estimator.